OpenStudy (zuha):

What is the value of this expression when x=5, y=7, and z=-8. X squared y+3z

OpenStudy (anonymous):

x^2 * y + 3z = 5^2 * 7 + 3(-8) = 151

OpenStudy (zuha):

Sooo, how did you solve it?

OpenStudy (anonymous):

Write the equation, and replace x with 5, y with 7, and z with -8

OpenStudy (zuha):

Thanks :)